OpenStudy (anonymous):

solve for y 5x-15y=20

OpenStudy (anonymous):

here's my answer \[y=\frac{ x }{ 3 } - \frac{ 4 }{ 3}\]

OpenStudy (anonymous):

but again it's none of the choices.

OpenStudy (astrophysics):

Hey lets take it slowly, so what was the first thing you did?

OpenStudy (anonymous):

ok hold on this equation thing is gonna take a while for me to use

OpenStudy (astrophysics):

You don't have to just tell me if you subtracted something or added what ever :)

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

\[5(x-3y)=20\]

OpenStudy (anonymous):

I factored out the common term which is 5

OpenStudy (anonymous):

tennisgril hold on

OpenStudy (astrophysics):

That's smart!!

OpenStudy (astrophysics):

But, we don't really need to do that here haha :)

OpenStudy (astrophysics):

\[5x-15y=20 \implies -15y=20-5x\] lets first subtract both sides by -5x

OpenStudy (anonymous):

after that i divided both sides by that same number

OpenStudy (astrophysics):

You can absolutely do that! But I think this maybe more clear for you and cause less confusion

OpenStudy (anonymous):

ok

OpenStudy (astrophysics):

Next we simply divide both sides by -15

OpenStudy (astrophysics):

\[y = \frac{ 20-5x }{ -15 } = -\frac{ 4 }{ 3 }+\frac{ x }{ 3 }\] so do you notice something compared to your answer?

OpenStudy (astrophysics):

Again it's exactly the same thing, remember i mentioned in the last post A+B = B+A just switch it around

OpenStudy (anonymous):

ok I dont see anything compared to that... :(

OpenStudy (astrophysics):

Are you sure? That is the answer, both of our answers are the same

OpenStudy (anonymous):

I know the thing is our answers don't match the answer choices given in the picture i attached

OpenStudy (anonymous):

i already did hold on lemme do it again

OpenStudy (astrophysics):

 it's right here

OpenStudy (astrophysics):

\[\frac{ x }{ 3 } = \frac{ 1 }{ 3 }x\]

OpenStudy (anonymous):

ok finally lemme see if it's correct

OpenStudy (anonymous):

yep! Thanks so much! If only there were more champions like you in this site

OpenStudy (astrophysics):

Yw :)